OpenStudy (jojokiw3):

Please explain the equation behind calculating this question. What is the variance of the following data?: 5, 6, 3, 6, 7, 2.

OpenStudy (s3a):

First find the mean (=average). mean = (5 + 6 + 3 + 6 + 7 + 2) / 6 = 29/6 =~ 4.83 Then, add each point minus the mean squared. That is, do the following.: (5 - 29/6)^2 + (6 - 29/6)^2 + (3 - 29/6)^2 + (6 - 29/6)^2 + (7 - 29/6)^2 + (2 - 29/6)^2 = 113/6 Then, divide the number 113/6 by one less than the amount of points you have. Since you have 6 points, divide by 5.: (113/6)/5 = 113/30 P.S. Don't be confused by the fractions, I just hate rounding numbers.

OpenStudy (error1603):

v